OpenStudy (anonymous):

find a polynomial function of degree 3 with -1,3,5 as zeros

OpenStudy (cwrw238):

(x + 3)(x - 3)(x - 5) expand this to get your polynomial

OpenStudy (cwrw238):

oops sorri the first 3 should be a 1 (x + 1)(x - 3)(x - 5)

OpenStudy (cwrw238):

can you expand this?

OpenStudy (anonymous):

no

OpenStudy (cwrw238):

do the last 2 factors first (x - 3)(x - 5) = x^2 -5x - 3x + 15 = x^2 - 8x + 15

OpenStudy (cwrw238):

now multiply by the (x + 1): (x + 1)(x^2 - 8x + 15) now multiply the 3 terms in second bracket by x then multiply them by + 1 and simplify the result i'll start you off: x*x^2 - 8x*x + 15*x +1*x^2 + do you follow the method?

OpenStudy (anonymous):

yes thank you

OpenStudy (cwrw238):

yw